Three packing crates of masses, M1 = 6 kg, M2 = 2 kg

and M3 = 8 kg are connected by a light string of

negligible mass that passes over the pulley as shown.

Masses M1 and M3 lies on a 30o incline plane which

slides down the plane. The coefficient of kinetic friction

on the incline plane is 0.28. Determine the tension in the string that connects m1 and m2


1
Expert's answer
2021-06-21T11:38:13-0400

"W_{1x}+ W_{3x} - fr = (m_1 + m_3) a\\

 N_1 - W_{1y} + N_3- W_{3y} = 0\\

 N_1+ N_3 = W_{1y} + W_{3y},"

"\sin 30° =\frac{ W_x}{W}\\

 W_x = W \sin 30°\\

 \cos 30° = \frac{W_{y} }{ W}\\

 W_{y} = W \cos 30°,"


"N_1+ N_3 = W_1 \cos 30° + W_3 \cos 30°\\

 W_{1x} + W_{3x} - μ (m_1 + m_3) g \cos30° = (m_1 + m_3) a\\

 a =\frac{ m_1g \sin 30° + m_3g \sin 30° - μ (m_1 + m_3) g \cos 30° }{m_1 + m_3}\\

 a = g \sin 30 °- μ g \cos30°,"


"a = 9.8 \sin 30 °- 0.28\cdot 9.8 \cos 30°\\

 a = 4.9 - 2.38\\

 a = 2.5 ~\frac{m}{s^2}."